# HiCoLoRA: Addressing Context-Prompt Misalignment via Hierarchical Collaborative LoRA for Zero-Shot DST

### v2

@@ Abstract
Zero-shot Dialog State Tracking (zs-DST) is essential for enabling Task-Oriented Dialog Systems (TODs) to generalize to new domains without costly data annotation. A central challenge lies in the semantic misalignment between dynamic dialog contexts and static prompts, leading to inflexible cross-layer coordination, domain interference, and catastrophic forgetting. To tackle this, we propose Hierarchical Collaborative Low-Rank Adaptation (HiCoLoRA), a framework that enhances zero-shot slot inference through robust prompt alignment. It features a hierarchical LoRA architecture for dynamic layer-specific processing (combining lower-layer heuristic grouping and higher-layer full interaction), integrates Spectral Joint Domain-Slot Clustering to identify transferable associations (feeding an Adaptive Linear Fusion Mechanism), and employs Semantic-Enhanced SVD Initialization (SemSVD-Init) to preserve pre-trained knowledge. Experiments on multi-domain datasets MultiWOZ and SGD show that HiCoLoRA outperforms baselines, achieving SOTA in zs-DST. Code is available at Github.
@@ 1 Introduction
Task-Oriented dialog Systems (TODs) help users complete specific tasks, such as restaurant reservations or taxi inquiries, through multi-turn natural language interactions Luo et al. (2024); Wang et al. (2024d). A core component enabling this functionality is dialog State Tracking (DST), which dynamically parses user inputs into structured slot-value pairs to infer intents and resolve ambiguities. However, zero-shot DST (zs-DST) faces a challenge: semantic misalignment between dynamic dialog contexts and static prompts, hindering adaptation to new domains.
To extend DST modules to unseen domains by leveraging existing knowledge and address data scarcity, zs-DST has emerged as a promising paradigm. While approaches include data augmentation He et al. (2025) and prompt engineering Liu et al. (2025b); Wang et al. (2024c); Aksu et al. (2023), parameter-efficient fine-tuning (PEFT), particularly Low-Rank Adaptation (LoRA) Wang et al. (2024a); Occhipinti et al. (2024), has gained prominence for zs-DST Yi et al. (2025); Aksu et al. (2023). LoRA freezes most pre-trained model parameters, updating only low-rank external matrices to enable efficient cross-domain generalization. Recent multi-LoRA variants, such as DualLoRA Luo et al. (2024), CoLA Zhou et al. (2025), HydraLoRA Tian et al. (2024), MTL-LoRA Yang et al. (2025), enhance adaptability through specialized adapters or cross-task collaboration. Despite these advances, structural limitations in context-prompt alignment persist, motivating our hierarchical approach.
@@
Despite these advancements, current LoRA based zs-DST methods face limitations. Data augmentation and prompt engineering approaches manipulate external data or rely on shallow input adjustments, which may not adequately cover a broad range of slot types or capture the nuanced complexities of dynamic dialog contexts. Similarly, PEFT methods rely on shallow input adjustments or local parameter modifications, which limits their adaptability to complex and dynamic dialog contexts. Specifically, a single LoRA projects features from different tasks in the same low-dimensional space. This can lead to intertask interference, hindering knowledge separation and limiting multitask adaptability. Although Multi-LoRA designs like DualLoRA Luo et al. (2024), CoLA Zhou et al. (2025), HydraLoRA Tian et al. (2024) and MTL-LoRA Yang et al. (2025) have attempted to mitigate these issues by introducing multipath adapters or exploring adaptive cross-task collaboration, limitations persist in practical applications. These limitations stem largely from a structural mismatch between dynamic dialog contexts and static prompts (as illustrated in Fig.1), manifesting itself in three critical research challenges: (RQ1) Rigid hierarchical designs hinder effective cross-layer weight sharing, limiting fine-grained semantic alignment in deeper layers. (RQ2) A single adaptation matrix conflates domain-agnostic and domain-specific signals, causing semantic confusion between domains. (RQ3) The use of random initialization for LoRA parameters can distort pre-trained knowledge and exacerbate catastrophic forgetting.
To address the three limitations, we propose Hierarchical Collaborative Low-Rank Adaptation (HiCoLoRA), a novel framework inspired by DualLoRA’s prompt augmentation Luo et al. (2024) and CoLA’s multi-LoRA grouping Zhou et al. (2025). Departing from “uniform layer processing", it introduces: (1) A Hierarchical Collaborative Architecture with lower-layer heuristic grouping, and higher-layer full interaction, resolving RQ1 via dynamic cross-layer coordination; (2) Spectral Joint Clustering and Adaptive Fusion disentangling domain-shared and specific semantics addressing RQ2; (3) Semantic-Enhanced SVD Initialization preserving pre-trained knowledge against RQ3.
Experiments on MultiWOZ and SGD datasets establish new SOTA results, achieving 5.4% and 9.4% JGA gains over DualLoRA, validating our framework’s success in fundamentally addressing RQ1-3 through hierarchical adaptation, semantic disentanglement and knowledge preservation, especially in high-overlap domains and sparse slots.
@@ 2 Related Work
zs-DST and Goal Accuracy. zs-DST is fundamental to TODs, with Joint Goal Accuracy (JGA) and Average Goal Accuracy (AGA) as key metric to evaluate performance. Early methods like TRADE Wu et al. (2019) and SUMBT Lee et al. (2019) laid the foundations for cross-domain generalization but relied on task-specific architectures. With pre-trained language models (PLMs), SimpleTOD Hosseini-Asl et al. (2020) reformulated DST as sequence generation, Yi et al. (2025) enhanced the few-shot capability through the enhancement of intent-driven dialog information. Recent advances on zs-DST include Prompter Aksu et al. (2023) with learnable prompts, DualLoRA Luo et al. (2024) with dual-path adapters and LUAS Wang et al. (2024d) with synthetic data, though these still face challenges in cross-layer semantic alignment and domain knowledge separation. HiCoLoRA fundamentally optimizes these challenges through its hierarchical cross-layer coordination and spectral domain-slot disentanglement,
Parameter-Efficient Fine-Tuning with LoRA. LoRA is an effective method for PEFT Zhang et al. (2025); Liu et al. (2025a); Jabbarvaziri & Lampe (2025); Zhang & Pilanci (2025); Wang et al. (2024b). For zero-shot scenarios, DualLoRA Luo et al. (2024) mitigates context-prompt misalignment via dual-path designs, while HydraLoRA Tian et al. (2024) uses MoE routers for subtask decoupling. Multitask adaptations such as CoLA Zhou et al. (2025) and MTL-LoRA Yang et al. (2025) enhance cross-task collaboration, and RoSA Nikdan et al. (2024) integrate routing/sparsity for efficiency. Initialization strategies such as PiSSA Meng et al. (2024), MiLoRA Zhang et al. (2024), improve knowledge preservation, but few address semantic alignment in dynamic dialog contexts. HiCoLoRA directly addresses this gap through hierarchical cross-layer semantic coordination and adaptive domain-slot disentanglement, enabling alignment in dynamic dialog contexts.
Layer-Specific Algorithms in Transformers. Xie et al. (2025); Wang et al. (2025); Liu et al. (2024); Du et al. (2020) all demonstrate that lower layers handle basic and detailed information, such as lexical semantics, grid features, and rapid computations, while higher layers focus on abstract and task-oriented processing, such as prediction, abstract planning, and semantic integration. Algorithm designs targeting this characteristic have improved task performance. Layer-specific algorithms also include Split Attention (partitioning attention across layers)Lin et al. (2025), Hierarchical LoRA (applying hierarchical LoRA patterns) Xiao et al. (2024); Guo et al. (2024), Dynamic Layer Replace (selective layer substitution)Xiong et al. (2024), and attention head pruning within layers He & Lin (2025); Zayed et al. (2024). These methods leverage the principle of unequal layer contributions across tasks, achieving computational or parameter reductions while improving metrics. Critically, they fail to resolve issues such as dynamic context-prompt misalignment induced by layer-specific adaptations and the lack of coordinated cross-layer interactions needed for semantic coherence, which challenges complex scenarios in zs-DST. HiCoLoRA thus aims to bridge this gap by introducing mechanisms for harmonizing hierarchical layer-wise adaptations and ensuring consistent cross-layer alignment essential for robust zs-DST performance.
@@ 3 Method
We propose Hierarchical Collaborative Low-Rank Adaptation (HiCoLoRA, Fig. 2), a method that enhances zero-shot slot inference in unseen domains through improved prompt alignment. HiCoLoRA employs a hierarchical architecture that moves beyond uniform layer-wise processing, dynamically integrating domain-agnostic (UniRep-LoRA) and domain-specific (SemAdapt-LoRA) semantics via adaptive fusion. Additionally, spectral clustering and SemSVD-Init optimize domain-slot representations to strengthen zero-shot generalization.
@@ 3.1 Universal Representation LoRA (UniRep-LoRA)
UniRep-LoRA is designed to efficiently capture domain-agnostic semantic information from the dialog context $\bm{x}_{ur}$, such as universal slots for time and location. By freezing the parameters of the pre-trained model $\bm{W}_{0}$ and updating only the low-rank matrices $\bm{B}_{ur}$ and $\bm{A}_{ur}$:
$\bm{h}_{ur}=\bm{W}_{0}x_{ur}+\bm{B}_{ur}\bm{A}_{ur}\bm{x}_{ur}.$ (1)
@@
Figure 2: Overview of the HiCoLoRA framework, which combines: (1) UniRep-LoRA and SemAdapt-LoRA with Adaptive Linear Fusion balancing domain-agnostic and domain-specific features; (2) Spectral Joint Domain-Slot Clustering disentangling domain semantics to guide fusion; (3) SemSVD-Init preserving pre-trained knowledge via singular value modulation. These synergistically address context-prompt misalignment, enhancing zero-shot slot inference.
@@ 3.2 Semantic Adaptation LoRA (SemAdapt-LoRA)
In contrast to UniRep-LoRA, which focuses on universal features, SemAdapt-LoRA is specifically tailored to optimize domain-specific prompts by dynamically adjusting their influence across different domains (relating to RQ2). To enable this domain-specific optimization, we introduce a Multi-Head Attention module: high-frequency dialog words from the train dataset serve as $\bm{Q}$, while slot descriptions function as $\bm{K}$ and $\bm{V}$, with the output denoted as $\bm{x}_{sa}$. Different attention heads allow for the capture of diverse semantic correlations between these two components. By improving the semantic alignment between high-frequency dialog information and slot descriptions, this setup provides SemAdapt-LoRA with more effective local semantic features as input, thus supporting its hierarchical collaborative mechanism to improve zero-shot slot inference performance. To achieve such fine-grained adaptive prompt processing, we further introduce two sets of trainable matrices: $\bm{A}_{sa}^{m}|_{m=1}^{M}$ for domain common prompt encoding and $\bm{B}_{sa}^{n}|_{n=1}^{N}$ for cluster-specific domain-slot reconstruction. The former compresses high-dimensional prompt semantics into a low-rank space, while the latter reconstructs low-rank features based on specific domain semantics, transforming general features into domain-specific representations. In detail, $M$ is the number of clusters of domains in domain clusters and $N$ is the number of clusters of domains.
@@
To ensure effective collaboration between $\bm{A}_{sa}^{m}|_{m=1}^{M}$ and $\bm{B}_{sa}^{n}|_{n=1}^{N}$, we propose a novel cross-layer collaborative module. Tailored to RQ1, it employs two interaction strategies across Transformer layers, aligned with their semantic roles: Lower layers encode Local Semantic Features serving as semantic atoms for higher layers, while higher layers model Global Semantic Features and guide lower-layer feature extraction via attention suppressing irrelevant associations. This paradigm transcends traditional Transformers’ uniform layer processing, forming a hierarchical semantic chain from local associations to global intent modeling.
To avoid inference latency with multi $\bm{A}_{sa}$ and $\bm{B}_{sa}$, we precompute their low-rank terms, merge them into model bias, and use these merged parameters during inference. like DualLoRA, it matches its efficiency despite more LORA matrices.
@@
Heuristic Grouping. For lower Transformer layers, tasked with encoding local semantic atoms, heuristic grouping is favored for its efficiency. It clusters semantically similar parameters to avoid irrelevant interactions, aggregating coherent local features that serve as building blocks for higher layers, laying a precise foundation for global processing.
$\bm{h}_{sa}=\bm{W}_{0}^{l}\bm{x}_{sa}+N\bm{B}_{sa}^{*}M\bm{A}_{sa}^{*}\bm{x}_{sa},$ (2)
@@
where $\bm{W}_{0}^{l}$ denotes lower-layer weights. The optimal algorithm for selecting matrices $\bm{A}_{sa}^{*}$ and $\bm{B}_{sa}^{*}$ is based on calculating the cosine similarity between the average vector of the slot clusters and the slot prompts. Specifically, $\bm{A}_{sa}^{*}$ refers to the matrix $\bm{A}$ corresponding to the category with the highest similarity between the embedding groups $\bm{x}_{sa}$ and the domain clusters $\mathcal{D}^{M}$, while $\bm{B}_{sa}^{*}$ denotes the matrix $\bm{B}$ associated with the category showing the highest similarity between $\bm{x}_{sa}$ and the slot prompt clusters $\mathcal{X}^{N}$. During training, differentiable selection is achieved via Gumbel-Softmax based on cluster similarity, while softmax is used to accelerate during inference.
Full Collaboration. Higher layers model global semantic connections through full collaboration, enabling comprehensive interactions between all encoded local atoms to capture implicit associations such as the link between train-arriveby and destination. This process, combined with attention-guided noise suppression from lower layers, resolves fine-grained alignment and enhances zero-shot generalization. The equation is shown below:
$\bm{h}_{sa}=\bm{W}_{0}^{h}\bm{x}_{sa}+\sum_{n=1}^{N}\bm{B}_{sa}^{n}\sum_{m=1}^{M}\bm{A}_{sa}^{m}\bm{x}_{sa},$ (3)
where $\bm{W}_{0}^{h}$ denotes high-layer weights.

### v3

@@ Abstract
Zero-shot Dialog State Tracking (zs-DST) is essential for enabling Task-Oriented Dialog Systems (TODs) to generalize to new domains without costly data annotation. A central challenge lies in the semantic misalignment between dynamic dialog contexts and static prompts, leading to inflexible cross-layer coordination, domain interference, and catastrophic forgetting. To tackle this, we propose Hierarchical Collaborative Low-Rank Adaptation (HiCoLoRA), a framework that enhances zero-shot slot inference through robust prompt alignment. It features a hierarchical LoRA architecture for dynamic layer-specific processing (combining lower-layer heuristic grouping and higher-layer full interaction), integrates Spectral Joint Domain-Slot Clustering to identify transferable associations (feeding an Adaptive Linear Fusion Mechanism), and employs Semantic-Enhanced SVD Initialization (SemSVD-Init) to preserve pre-trained knowledge. Experiments on multi-domain datasets MultiWOZ and SGD show that HiCoLoRA outperforms baselines, achieving SOTA in zs-DST. Code is available at Anonymous Github.
HiCoLoRA: Addressing Context-Prompt Misalignment via Hierarchical Collaborative LoRA for Zero-Shot DST
Shuyu Zhang1thanks: These authors contributed equally to this work.,  Yifan Wei2 *,  Xinru Wang3,  Yanmin Zhu1thanks: Corresponding Author.,   Yangfan He4,  Yixuan Weng5, Bin Li6 †Yujie Liu7 1Department of Computer Science and Engineering, Shanghai Jiao Tong University 2School of Computer Science and Engineering, Beihang University 3Data Analytics for Business, University of Sydney 4University of Minnesota – Twin Cities 5Westlake University 6Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences 7School of Information Engineering, Beijing Institute of Graphic Communication carsonz@sjtu.edu.cn, weiyifan@buaa.edu.cn, xinruwang1999@outlook.com, yzhu@cs.sjtu.edu.cn, he000577@umn.edu, wengsyx@gmail.com, b.li2@siat.ac.cn, 18764008899@163.com
@@ 1 Introduction
Task-Oriented Dialog Systems (TODs) assist users to complete specific tasks, such as hotel or flight bookings Luo et al. (2024); Wang et al. (2024d). To generalize to new domains without extensive training, TODs rely on zero-shot Dialog State Tracking (zs-DST), yet face challenges of data scarcity and domain adaptation. A central obstacle is the semantic misalignment between dynamic dialog contexts and static prompts. This misalignment limits effective cross domain transfer.
To address data scarcity and enable generalization, existing approaches for zs-DST include data augmentation He et al. (2025), prompt engineering Liu et al. (2025b); Wang et al. (2024c); Aksu et al. (2023), and parameter-efficient fine-tuning (PEFT). Among PEFT methods, Low-Rank Adaptation (LoRA) Wang et al. (2024a); Occhipinti et al. (2024) has gained prominence for zs-DST Yi et al. (2025); Aksu et al. (2023) due to its ability to adapt large models with minimal parameters. Building upon LoRA, recent multi-LoRA variants such as DualLoRA Luo et al. (2024), CoLA Zhou et al. (2025), HydraLoRA Tian et al. (2024), and MTL-LoRA Yang et al. (2025) enhance adaptability through specialized adapters or cross-task collaboration. The semantic misalignment arises from the inherent layer-wise processing in Transformers: lower layers capture local semantic atoms while higher layers integrate them into global intent representations Liu et al. (2024). Despite these advances, existing LoRA variants treat all layers uniformly, failing to coordinate these distinct semantic roles, especially in zero-shot scenarios where domain-specific cues are scarce.
These limitations stem from a structural mismatch between dynamic dialog contexts and static prompts as illustrated in Fig. 1, which manifests in three critical research questions: (RQ1) Rigid hierarchical designs hinder effective cross-layer weight sharing, limiting fine-grained semantic alignment in deeper layers. (RQ2) A single adaptation matrix conflates domain-agnostic and domain specific signals, causing semantic confusion between domains. (RQ3) The use of random initialization for LoRA parameters can distort pre-trained knowledge and exacerbate catastrophic forgetting.
To address the three limitations, we propose Hierarchical Collaborative Low-Rank Adaptation (HiCoLoRA), a novel framework inspired by DualLoRA’s prompt augmentation Luo et al. (2024) and CoLA’s multi-LoRA grouping Zhou et al. (2025). Departing from “uniform layer processing”, our contributions are: (1) A Hierarchical Collaborative Architecture with lower-layer heuristic grouping and higher-layer full interaction, resolving RQ1 via dynamic cross-layer coordination; (2) Spectral Joint Clustering and Adaptive Fusion disentangling domain-shared and specific semantics addressing RQ2; (3) Semantic-Enhanced SVD Initialization preserving pre-trained knowledge against RQ3.
Extensive experiments on MultiWOZ and SGD multi-domain datasets demonstrate that HiCoLoRA significantly outperforms previous SOTA methods, validating the effectiveness of our proposed framework in addressing the core challenges of zs-DST.
@@ 2 Related Work
Layer-Specific Algorithms in Transformers. The hierarchical processing in Transformers is well established: lower layers capture local semantic atoms, while higher layers integrate them into global intent representations Liu et al. (2024); Wang et al. (2025). To exploit this asymmetry, various layer specific algorithms have been proposed, including hierarchical LoRA Xiao et al. (2024); Guo et al. (2024), dynamic layer replacement Xiong et al. (2024), attention head pruning He and Lin (2025); Zayed et al. (2024), and split attention mechanisms Lin et al. (2025). While these methods improve efficiency by leveraging unequal layer contributions, they typically treat layers uniformly, lacking mechanisms to coordinate their distinct semantic roles, particularly crucial for dynamic dialog contexts where fine grained alignment between evolving utterances and static prompts is required. HiCoLoRA addresses this gap through explicit hierarchical collaboration, enabling dynamic cross-layer coordination.
PEFT with LoRA. LoRA and its variants have become prominent for zs-DST due to their parameter efficiency Zhang et al. (2025); Liu et al. (2025a); Jabbarvaziri and Lampe (2025). DualLoRA Luo et al. (2024) uses dual adapters to align contexts and prompts, while multi-adapter approaches like HydraLoRA Tian et al. (2024), CoLA Zhou et al. (2025), and MTL-LoRA Yang et al. (2025) enhance cross-task collaboration. Methods like RoSA Nikdan et al. (2024) combine low-rank and sparse adaptations for efficiency, while initialization strategies like PiSSA Meng et al. (2024) and MiLoRA Zhang et al. (2024) aim to better preserve pre-trained knowledge. Spectral based adaptations have also been explored Zhang and Pilanci (2025); Wang et al. (2024b). Yet, these approaches still struggle with the semantic misalignment between dynamic dialog contexts and static slot prompts, often conflating domain agnostic and domain specific signals, which hinders zero-shot generalization. HiCoLoRA directly tackles this via spectral joint clustering and adaptive fusion, enabling disentangled and aligned representations.
@@
zs-DST and Goal Accuracy. zs-DST aims to generalize to unseen domains without annotated data. Early methods like TRADE Wu et al. (2019) and SUMBT Lee et al. (2019) relied on task-specific architectures. With the advent of PLMs, generation based approaches like SimpleTOD Hosseini-Asl et al. (2020) and intent enhanced methods Yi et al. (2025) have advanced the field. Prompt based methods like Prompter Aksu et al. (2023) and dual adapter approaches like DualLoRA Luo et al. (2024) further improve cross-domain transfer, while synthetic data methods like LUAS Wang et al. (2024d) address data scarcity. However, these approaches remain limited by rigid layer processing, insufficient semantic disentanglement, and knowledge distortion during adaptation. HiCoLoRA fundamentally optimizes these challenges through its hierarchical cross-layer coordination and spectral domain-slot disentanglement.
@@ 3 Method
We propose Hierarchical Collaborative Low-Rank Adaptation (HiCoLoRA, Fig. 2) to address the research questions: A hierarchical collaborative architecture resolves cross-layer rigidity (RQ1) by enabling dynamic layer specific coordination; Spectral joint domain-slot clustering that disentangles domain shared and domain specific semantics (RQ2), guiding adaptive fusion between domain-agnostic (UniRep-LoRA) and domain specific (SemAdapt-LoRA) representations; and Semantic Enhanced SVD Initialization (SemSVD-Init) that preserves pre-trained knowledge against catastrophic forgetting (RQ3). These components systematically mitigate context prompt misalignment and strengthen zero-shot generalization.
@@ 3.1 Universal Representation LoRA (UniRep-LoRA)
UniRep-LoRA is designed to efficiently capture domain-agnostic semantic information from the dialog context $\boldsymbol{x}_{ur}$, such as universal slots for time and location. By freezing the parameters of the pre-trained model $\boldsymbol{W}_{0}$ and updating only the low-rank matrices $\boldsymbol{B}_{ur}$ and $\boldsymbol{A}_{ur}$:
$\boldsymbol{h}_{ur}=\boldsymbol{W}_{0}x_{ur}+\boldsymbol{B}_{ur}\boldsymbol{A}_{ur}\boldsymbol{x}_{ur}.$ (1)
@@
Figure 2: The HiCoLoRA framework combines: (1) UniRep-LoRA and SemAdapt-LoRA with Adaptive Linear Fusion balancing domain-agnostic and domain-specific features; (2) Spectral Joint Domain-Slot Clustering disentangling domain semantics to guide fusion; (3) SemSVD-Init preserving pre-trained knowledge via singular value modulation. These synergistically address context-prompt misalignment, enhancing zero-shot slot inference.
@@ 3.2 Semantic Adaptation LoRA (SemAdapt-LoRA)
Unlike UniRep-LoRA which captures universal features, SemAdapt-LoRA specializes in domain specific prompt optimization, dynamically adjusting their influence across domains (addressing RQ2). We first employ a Multi-Head Attention module to enhance semantic alignment between dialog contexts and slot descriptions: high frequency dialog words from the training set serve as $\boldsymbol{Q}$, while slot descriptions act as $\boldsymbol{K}$ and $\boldsymbol{V}$, yielding the aligned representation $\boldsymbol{x}_{sa}$. This allows different heads to capture diverse semantic correlations, providing richer local semantic inputs for subsequent hierarchical processing.
To enable fine grained adaptation, we introduce two sets of trainable low-rank matrices: $\boldsymbol{A}_{sa}^{m}|_{m=1}^{M}$ for domain common prompt encoding, and $\boldsymbol{B}_{sa}^{n}|_{n=1}^{N}$ for cluster specific domain-slot reconstruction. Here, $M$ denotes the number of identified domain clusters, and $N$ the number of semantic clusters for slot prompts. $\boldsymbol{A}_{sa}^{m}$ compresses high dimensional prompt semantics into a low-rank space, while $\boldsymbol{B}_{sa}^{n}$ reconstructs domain specific representations from these compressed features.
@@
To ensure effective collaboration between $\boldsymbol{A}_{sa}^{m}$ and $\boldsymbol{B}_{sa}^{n}$ across layers (addressing RQ1), we design a cross-layer collaborative module that respects the distinct semantic roles of Transformer layers: lower layers capture local semantic features serving as semantic atoms, while higher layers model global semantic features and guide lower-layer feature extraction via attention based suppression of irrelevant associations. This hierarchical approach moves beyond uniform layer processing, forming a semantic chain from local cues to global intent.
@@
Heuristic Grouping. For lower layers that encode local semantic atoms, we adopt heuristic grouping to efficiently aggregate semantically similar parameters, reducing irrelevant interactions and providing a precise foundation for higher layer processing:
$\boldsymbol{h}_{sa}=\boldsymbol{W}_{0}^{l}\boldsymbol{x}_{sa}+N\boldsymbol{B}_{sa}^{*}M\boldsymbol{A}_{sa}^{*}\boldsymbol{x}_{sa},$ (2)
@@
where $\boldsymbol{W}_{0}^{l}$ denotes lower-layer weights. Matrices $\boldsymbol{A}_{sa}^{*}$ and $\boldsymbol{B}_{sa}^{*}$ are selected based on cosine similarity between the input embeddings $\boldsymbol{x}_{sa}$ and the cluster centroids of domains ($\mathcal{D}^{M}$) and slot prompts ($\mathcal{X}^{N}$), respectively. During training, differentiable selection is performed via Gumbel-Softmax, while at inference, we use softmax for efficiency.
Full Collaboration. Higher layers integrate all local semantic atoms through full collaboration, capturing implicit cross slot associations, such as train-arriveby and destination. This process can suppress irrelevant signals via attention guidance from lower layers:
$\boldsymbol{h}_{sa}=\boldsymbol{W}_{0}^{h}\boldsymbol{x}_{sa}+\sum_{n=1}^{N}\boldsymbol{B}_{sa}^{n}\sum_{m=1}^{M}\boldsymbol{A}_{sa}^{m}\boldsymbol{x}_{sa},$ (3)
where $\boldsymbol{W}_{0}^{h}$ denotes higher-layer weights.